# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v4

@@ 2.4 Dislocation strain field evolution
Based on observations of dislocation pileup evolution upon introduction of hydrogen, it has been previously inferred that hydrogen reduces the interaction between dislocations by changing their stress fields [19, 44]. To quantify the changes in the strain field surrounding the large dislocation, we compare slices though the 3D reconstructed volume where $\mathbf{b}$ and $\bar{1}11]$ (the normal to the initial glide plane of the large dislocation) are in plane and $\xi$ is normal to the slice (i.e. where the large dislocation has predominantly edge character). These slices are shown in Figure 5a, b. A theoretical model, devoid of hydrogen, of the edge dislocation strain field projected along $[111]$ is shown in Figure 5c (Experimental Section), and compared to the experimental strain field by considering circular line profiles drawn at a 30 nm radius from the dislocation core (Figure 5d). Agreement between the theoretical model and the experiment is initially good, but worsens as the hydrogen concentration increases.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e,f. Figure 5g shows how the maximum and minimum values, averaged over $\pm\pi/4$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon the introduction of hydrogen [67].
At the nanoscale, we attribute this elastic strain reduction to the formation of a hydrogen atmosphere in the vicinity of the dislocation. Dislocations are relatively large and strong hydrogen trapping sites ($\approx 27\ \mathrm{kJ\ mol^{-1}}$) [25], so the hydrogen atoms surrounding the dislocations are relatively immobile at room temperature [27]. Therefore, the evolution of this localized accumulation can be probed using the heterogeneous strain. Figure 5g also shows that the maximum $\varepsilon_{111}$ deviates more from the elastic hydrogen-free model than the minimum $\varepsilon_{111}$. This suggests that hydrogen prefers to reside in the tensile region of the dislocation [19]. Alternatively, the reduction in tensile strain could also be due to the hydrogen-facilitated diffusion of solutes, resulting in solute segregation [56], but this is unlikely since we are at room temperature.
Our observations reveal hydrogen elastic shielding (Figure 5) which has been predicted by simulations [19], and indirectly inferred from in situ TEM observations of dislocation pileup evolution following the introduction of hydrogen [44]. These direct BCDI measurements of subtle changes in dislocation strain field are important, since they provide fundamental validation of hydrogen-induced dislocation stress shielding, which is core to the HELP mechanism.
@@ 2.5 Proposed mechanisms
We now consider whether dislocation mobility can be introduced by internal stresses induced by hydrogen concentration gradients caused by cathodic hydrogen charging [68]. As hydrogen diffuses into the SS disk from the top, lattice swelling ensues [54, 29], creating a hydrostatic internal stress. There is an in-plane boundary constraint from the uncharged bulk that generates an in-plane stress on the grain, such that the in-plane strain remains zero, as required for compatibility [69]. We have modeled the experimental conditions and predicted the hydrogen concentration as a function of time and depth below the surface of the disk (Supporting Information and Figure LABEL:supp-fig:H_concentration_profilea, Supporting Information). The resulting strain distributions can be compared to the experimental data. In Fig. LABEL:supp-fig:Strain_comparison, Supporting Information, we observe that there is an additional background strain field in the measured $\varepsilon_{111}+e_{111}$ data that is not directly captured by the hydrogen-induced strain gradient combined with the large dislocation strain field model, $\varepsilon_{111,\ \mathrm{model}}+\varepsilon_{111,\ \mathrm{b.c.}}$. We do note that this background remains largely unchanged throughout the charging, suggesting that stresses due to neighboring grains remain largely unchanged during hydrogen uptake.
@@
The hydrogen uptake creates a resolved shear stress that acts on the dislocation. The in-plane stress has been applied in the DDD model, which shows only a slight lift in part of the dislocation toward the unpinning direction (orange arrows in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information). Computation of the resulting Peach–Koehler (PK) force on the large dislocation reveals a significant component in the out-of-plane direction of dislocation climb (purple arrows in Figure LABEL:supp-fig:Fig4_bc, Supporting Information). Interestingly, the PK force also has an in-plane component opposite the direction of glide, which explains why the DDD model does not exhibit gliding behavior in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information. This allows us to conclude that, while hydrogen gradient-induced stresses are significant, they do not appear to be the primary factor controlling dislocation glide. Instead, the hydrogen stress shielding effect [44] reduces the large dislocation’s effective Burgers vector, lowering the PK force and creating an energetically favorable condition where the dislocation line straightens, as we observe experimentally.
@@
Our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms. A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by DFT studies on pure metals using the Peierls–Nabarro model [70]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [71] and reduce the dislocation strain field [67], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [19, 16, 44] may no longer hold.
@@ 3 Conclusion
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as third-generation synchrotrons are being upgraded to fourth-generation facilities, offering orders of magnitude greater coherent flux for BCDI experiments that will enhance spatial and temporal resolution. This could enable exciting possibilities for exploring hydrogen-induced evolution in adjacent grains or extended grains using Bragg ptychography [72]. In situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Using the example of 316 SS, our results show the hydrogen-driven evolution of a single dislocation undergoing glide, followed by climb. We propose that glide is initiated by the diffusion of hydrogen into the lattice, where it reduces pinning strength and thus enables the dislocation to glide in-plane, reducing its line length and thus elastic energy. Our strain measurements also show that hydrogen shields the elastic strain field associated with the dislocation. With increasing hydrogen concentration, we observe rapid dislocation climb that is not expected at room temperature. This suggests that hydrogen alters the energy landscape, such that vacancies and interstitials become more readily available at the dislocation line, thus enabling climb motion. Understanding these internal mechanisms is key to designing alloys with enhanced resistance to hydrogen-induced degradation, such as engineering micro- and nano-structures to mitigate dislocation movement or vacancy formation [73]. These hydrogen-resistant materials are urgently needed to enable the green hydrogen economy to help achieve decarbonization targets.
@@ 4.1 Sample preparation
Powder X-ray diffraction (Figure LABEL:supp-fig:powder_XRD, Supporting Information) showed that after HPT, the sample was predominately composed of deformation-induced martensite. The heat treatment removed macroscale residual stresses and the sample underwent recovery and recrystallization to form austenite [74, 75]. The measured austenite grain was likely formed during heat treatment due to the relatively low misorientation and very low observed strain (relative to the average lattice parameter of the grain) near the surface of the grain. It was noted that carbides could have formed during the annealing, which could attract solutes that have become mobile due to hydrogen.
@@ 4.2 BCDI measurements
BCDI relies on inverting far-field, oversampled [76], 3D coherent X-ray diffraction patterns (CXDPs) for a particular hklhkl reflection from a finite crystalline material. This was done using phase retrieval algorithms [52] to obtain the grain morphology and associated phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$. Dislocation line positions can be readily identified from the positions of phase vortex loci [46], thus making BCDI a valuable tool to study in situ or in operando defect evolution [46, 47, 48, 49, 77]. The strain fields associated with the dislocations can be obtained from the phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$, since $\mathbf{\psi}_{hkl}(\mathbf{r})$ is a projection of the lattice displacement field, $\mathbf{u}_{hkl}(\mathbf{r})$, onto the scattering vector, $\mathbf{Q}_{hkl}$ [78].
BCDI was performed at beamline 34-ID-C at the Advanced Photon Source (APS) at Argonne National Laboratory, USA. An in situ confocal microscope was used to position the X-ray beam within the disk. The $111$ Bragg peak ($2\theta=34.7$°) was measured for multiple grains to screen for the best candidate grain. Although each candidate’s position perpendicular to the X-ray beam could be easily aligned, the position of the candidate along the beam needs to be determined to ensure the grain is on the axis of rotation. An approach presented by Shabalin et al. [79] was used to accomplish this.
@@ 4.4 Phase retrieval
The reconstructions were processed using a MATLAB phase retrieval package [46]. The reconstructions for each time point were seeded using a reconstruction from eight repeated CXDPs (combined following the same procedure) of the grain before the charging solution was pumped into the flow cell. This seed was reconstructed starting with a random guess. A guided phasing approach [80] with 100 individuals and four generations was used with a geometric average breeding mode and a low to high-resolution scheme [81]. For each generation and population, a block of 20 error reduction (ER) and 180 hybrid input-output (HIO) iterations, with $\beta=0.9$, was repeated three times. This was followed by 20 ER iterations to return the final object. The shrinkwrap algorithm [82] with a threshold of 0.08 was used to update the real-space support every iteration. The best reconstruction was determined using a sharpness criterion, appropriate for crystals containing defects [83]. The seed reconstruction was then used as the initial guess for the reconstructions at all time points, which followed the same guided phase retrieval procedure as the seed.
@@ 4.6 Hydrogen concentration
The volumetric strain, $\varepsilon_{\mathrm{vol}}$, associated with the swelling of the grain due to hydrogen uptake can be written as [69],
@@
where $C^{(A)}$ is the defect concentration for defect type $A$, and $\Omega_{R}^{(A)}$ is the relaxation volume for the specific defect type. The FCC metal lattice has two interstitial sites available for accommodating hydrogen atoms: octahedral (O) and tetrahedral (T) sites. In transition metals with FCC lattice, dissolved hydrogen atoms preferentially occupy the O site with larger free space than the T site [29]. $\Omega_{R}=0.200\pm 0.005$ was used for interstitial hydrogen in austenitic stainless steels [54]. For the calculation, this value was assumed to be similar for FCC 316 SS, and that all the interstitial hydrogen atoms reside in the O sites.
@@
The hydrogen concentrations presented in Figure 1b are reasonable when compared to values obtained from charging hydrogen into 316L $S$ [84].
@@ 4.8 Dislocation position identification
The dislocation line is identified as the spatial positions of the locus of the phase vortices, referred to as nodes. Each dislocation was composed of many nodes, joined by edges based on the MATLAB graph object. The nodes were determined automatically by integrating the derivatives of the complex exponential of the phase $(e^{i\psi_{hkl}(\mathbf{r})})$ [85] and then selecting the maximum value as the dislocation node position if the value exceeded $0.75\pi$. The dislocation lines were found not to overlap perfectly in the reconstructions, which were attributed to noise and the limited spatial resolution. To increase overlap, the reconstructions were subpixel-shifted [86], on the order of a few pixels, or up to $\approx 30$ nm. The dislocation lines terminated on the surface of the grain.
@@ 4.9.1 Overview
DDD was used to simulate the motion of the dislocation structure and compare to the experimental result. The DDD simulations here are nodal, based on discrete straight segments [87, 88] presented by Yu et al. [89]. Dislocation motion is calculated by evaluating the velocity $\textbf{V}_{k}$ of node $k$ at position $\textbf{X}_{k}$ through a mobility law (shown in the following section), which describes mobility in various modes like glide, climb, and cross-slip.

### v5

@@ 2.4 Dislocation strain field evolution
Based on observations of dislocation pileup evolution upon introduction of hydrogen, it has been previously inferred that hydrogen reduces the interaction between dislocations by changing their stress fields [Birnbaum1994, Ferreira1998]. To quantify the changes in the strain field surrounding the large dislocation, we compare slices though the 3D reconstructed volume where $\mathbf{b}$ and $\bar{1}11]$ (the normal to the initial glide plane of the large dislocation) are in plane and $\xi$ is normal to the slice (i.e. where the large dislocation has predominantly edge character). These slices are shown in Figure 5a, b. A theoretical model, devoid of hydrogen, of the edge dislocation strain field projected along $[111]$ is shown in Figure 5c (Experimental Section), and compared to the experimental strain field by considering circular line profiles drawn at a 30 nm radius from the dislocation core (Figure 5d). Agreement between the theoretical model and the experiment is initially good, but worsens as the hydrogen concentration increases.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e,f. Figure 5g shows how the maximum and minimum values, averaged over $\pm\pi/4$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon the introduction of hydrogen [Sofronis1995].
At the nanoscale, we attribute this elastic strain reduction to the formation of a hydrogen atmosphere in the vicinity of the dislocation. Dislocations are relatively large and strong hydrogen trapping sites ($\approx 27\ \mathrm{kJ\ mol^{-1}}$) [Choo1982a], so the hydrogen atoms surrounding the dislocations are relatively immobile at room temperature [Myers1992]. Therefore, the evolution of this localized accumulation can be probed using the heterogeneous strain. Figure 5g also shows that the maximum $\varepsilon_{111}$ deviates more from the elastic hydrogen-free model than the minimum $\varepsilon_{111}$. This suggests that hydrogen prefers to reside in the tensile region of the dislocation [Birnbaum1994]. Alternatively, the reduction in tensile strain could also be due to the hydrogen-facilitated diffusion of solutes, resulting in solute segregation [Kirchheim2007a], but this is unlikely since we are at room temperature.
Our observations reveal hydrogen elastic shielding (Figure 5) which has been predicted by simulations [Birnbaum1994], and indirectly inferred from in situ TEM observations of dislocation pileup evolution following the introduction of hydrogen [Ferreira1998]. These direct BCDI measurements of subtle changes in dislocation strain field are important, since they provide fundamental validation of hydrogen-induced dislocation stress shielding, which is core to the HELP mechanism.
@@ 2.5 Proposed mechanisms
We now consider whether dislocation mobility can be introduced by internal stresses induced by hydrogen concentration gradients caused by cathodic hydrogen charging [Rozenak1983]. As hydrogen diffuses into the SS disk from the top, lattice swelling ensues [Ulmer1993, Fukai2005], creating a hydrostatic internal stress. There is an in-plane boundary constraint from the uncharged bulk that generates an in-plane stress on the grain, such that the in-plane strain remains zero, as required for compatibility [Hofmann2015]. We have modeled the experimental conditions and predicted the hydrogen concentration as a function of time and depth below the surface of the disk (Supporting Information and Figure LABEL:supp-fig:H_concentration_profilea, Supporting Information). The resulting strain distributions can be compared to the experimental data. In Fig. LABEL:supp-fig:Strain_comparison, Supporting Information, we observe that there is an additional background strain field in the measured $\varepsilon_{111}+e_{111}$ data that is not directly captured by the hydrogen-induced strain gradient combined with the large dislocation strain field model, $\varepsilon_{111,\ \mathrm{model}}+\varepsilon_{111,\ \mathrm{b.c.}}$. We do note that this background remains largely unchanged throughout the charging, suggesting that stresses due to neighboring grains remain largely unchanged during hydrogen uptake.
@@
The hydrogen uptake creates a resolved shear stress that acts on the dislocation. The in-plane stress has been applied in the DDD model, which shows only a slight lift in part of the dislocation toward the unpinning direction (orange arrows in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information). Computation of the resulting Peach–Koehler (PK) force on the large dislocation reveals a significant component in the out-of-plane direction of dislocation climb (purple arrows in Figure LABEL:supp-fig:Fig4_bc, Supporting Information). Interestingly, the PK force also has an in-plane component opposite the direction of glide, which explains why the DDD model does not exhibit gliding behavior in Figure LABEL:supp-fig:dislocation_dynamics_models_stress, Supporting Information. This allows us to conclude that, while hydrogen gradient-induced stresses are significant, they do not appear to be the primary factor controlling dislocation glide. Instead, the hydrogen stress shielding effect [Ferreira1998] reduces the large dislocation’s effective Burgers vector, lowering the PK force and creating an energetically favorable condition where the dislocation line straightens, as we observe experimentally.
@@
Our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms. A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by DFT studies on pure metals using the Peierls–Nabarro model [Kumar2021]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [Yu2020] and reduce the dislocation strain field [Sofronis1995], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [Birnbaum1994, Robertson2015, Ferreira1998] may no longer hold.
@@ 3 Conclusion
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as third-generation synchrotrons are being upgraded to fourth-generation facilities, offering orders of magnitude greater coherent flux for BCDI experiments that will enhance spatial and temporal resolution. This could enable exciting possibilities for exploring hydrogen-induced evolution in adjacent grains or extended grains using Bragg ptychography [Godard2011]. In situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Using the example of 316 SS, our results show the hydrogen-driven evolution of a single dislocation undergoing glide, followed by climb. We propose that glide is initiated by the diffusion of hydrogen into the lattice, where it reduces pinning strength and thus enables the dislocation to glide in-plane, reducing its line length and thus elastic energy. Our strain measurements also show that hydrogen shields the elastic strain field associated with the dislocation. With increasing hydrogen concentration, we observe rapid dislocation climb that is not expected at room temperature. This suggests that hydrogen alters the energy landscape, such that vacancies and interstitials become more readily available at the dislocation line, thus enabling climb motion. Understanding these internal mechanisms is key to designing alloys with enhanced resistance to hydrogen-induced degradation, such as engineering micro- and nano-structures to mitigate dislocation movement or vacancy formation [Li2020c]. These hydrogen-resistant materials are urgently needed to enable the green hydrogen economy to help achieve decarbonization targets.
@@ 4.1 Sample preparation
Powder X-ray diffraction (Figure LABEL:supp-fig:powder_XRD, Supporting Information) showed that after HPT, the sample was predominately composed of deformation-induced martensite. The heat treatment removed macroscale residual stresses and the sample underwent recovery and recrystallization to form austenite [Bartova2022, Naghizadeh2018]. The measured austenite grain was likely formed during heat treatment due to the relatively low misorientation and very low observed strain (relative to the average lattice parameter of the grain) near the surface of the grain. It was noted that carbides could have formed during the annealing, which could attract solutes that have become mobile due to hydrogen.
@@ 4.2 BCDI measurements
BCDI relies on inverting far-field, oversampled [Sayre1952], 3D coherent X-ray diffraction patterns (CXDPs) for a particular hklhkl reflection from a finite crystalline material. This was done using phase retrieval algorithms [Fienup1982] to obtain the grain morphology and associated phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$. Dislocation line positions can be readily identified from the positions of phase vortex loci [Clark2015], thus making BCDI a valuable tool to study in situ or in operando defect evolution [Clark2015, Orr2023, Singer2018, Yau2017b, Gorobtsov2023]. The strain fields associated with the dislocations can be obtained from the phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$, since $\mathbf{\psi}_{hkl}(\mathbf{r})$ is a projection of the lattice displacement field, $\mathbf{u}_{hkl}(\mathbf{r})$, onto the scattering vector, $\mathbf{Q}_{hkl}$ [Robinson2009].
BCDI was performed at beamline 34-ID-C at the Advanced Photon Source (APS) at Argonne National Laboratory, USA. An in situ confocal microscope was used to position the X-ray beam within the disk. The $111$ Bragg peak ($2\theta=34.7$°) was measured for multiple grains to screen for the best candidate grain. Although each candidate’s position perpendicular to the X-ray beam could be easily aligned, the position of the candidate along the beam needs to be determined to ensure the grain is on the axis of rotation. An approach presented by Shabalin et al. [Shabalin2023] was used to accomplish this.
@@ 4.4 Phase retrieval
The reconstructions were processed using a MATLAB phase retrieval package [Clark2015]. The reconstructions for each time point were seeded using a reconstruction from eight repeated CXDPs (combined following the same procedure) of the grain before the charging solution was pumped into the flow cell. This seed was reconstructed starting with a random guess. A guided phasing approach [Chen2007] with 100 individuals and four generations was used with a geometric average breeding mode and a low to high-resolution scheme [McCallum1989]. For each generation and population, a block of 20 error reduction (ER) and 180 hybrid input-output (HIO) iterations, with $\beta=0.9$, was repeated three times. This was followed by 20 ER iterations to return the final object. The shrinkwrap algorithm [Marchesini2003] with a threshold of 0.08 was used to update the real-space support every iteration. The best reconstruction was determined using a sharpness criterion, appropriate for crystals containing defects [Ulvestad2017b]. The seed reconstruction was then used as the initial guess for the reconstructions at all time points, which followed the same guided phase retrieval procedure as the seed.
@@ 4.6 Hydrogen concentration
The volumetric strain, $\varepsilon_{\mathrm{vol}}$, associated with the swelling of the grain due to hydrogen uptake can be written as [Hofmann2015],
@@
where $C^{(A)}$ is the defect concentration for defect type $A$, and $\Omega_{R}^{(A)}$ is the relaxation volume for the specific defect type. The FCC metal lattice has two interstitial sites available for accommodating hydrogen atoms: octahedral (O) and tetrahedral (T) sites. In transition metals with FCC lattice, dissolved hydrogen atoms preferentially occupy the O site with larger free space than the T site [Fukai2005]. $\Omega_{R}=0.200\pm 0.005$ was used for interstitial hydrogen in austenitic stainless steels [Ulmer1993]. For the calculation, this value was assumed to be similar for FCC 316 SS, and that all the interstitial hydrogen atoms reside in the O sites.
@@
The hydrogen concentrations presented in Figure 1b are reasonable when compared to values obtained from charging hydrogen into 316L SS [Brass2006].
@@ 4.8 Dislocation position identification
The dislocation line is identified as the spatial positions of the locus of the phase vortices, referred to as nodes. Each dislocation was composed of many nodes, joined by edges based on the MATLAB graph object. The nodes were determined automatically by integrating the derivatives of the complex exponential of the phase $(e^{i\psi_{hkl}(\mathbf{r})})$ [Yang2022c] and then selecting the maximum value as the dislocation node position if the value exceeded $0.75\pi$. The dislocation lines were found not to overlap perfectly in the reconstructions, which were attributed to noise and the limited spatial resolution. To increase overlap, the reconstructions were subpixel-shifted [Guizar-Sicairos2008], on the order of a few pixels, or up to $\approx 30$ nm. The dislocation lines terminated on the surface of the grain.
@@ 4.9.1 Overview
DDD was used to simulate the motion of the dislocation structure and compare to the experimental result. The DDD simulations here are nodal, based on discrete straight segments [Arsenlis:MSMSE:2007, Cai:MSEA:2004] presented by Yu et al. [Yu:JotMaPoS:2019]. Dislocation motion is calculated by evaluating the velocity $\textbf{V}_{k}$ of node $k$ at position $\textbf{X}_{k}$ through a mobility law (shown in the following section), which describes mobility in various modes like glide, climb, and cross-slip.